OpenStudy (anonymous):

Find S11 for 3 + 1 + (-1) + (-3) +… A. -77 B. -75 C. -73 D. -71

OpenStudy (jdoe0001):

hmm doesn't look like a sequence to me for one

OpenStudy (jdoe0001):

can you post a quick screenshot of the material? so we can see what you mean

OpenStudy (astrophysics):

Nope, it's a series

OpenStudy (jdoe0001):

hmm ok

OpenStudy (astrophysics):

So we can use sum of an arithmetic series equation I guess which is \[S_n = \frac{ n }{ 2 }[2a+(n-a)d]\] a = first term in series, d = common difference

OpenStudy (astrophysics):

n is what you're looking for in this case n = 11

OpenStudy (astrophysics):

So what's the common difference, look at your previous post what I wrote when you left

OpenStudy (astrophysics):

So I'll post it again for reference, the common difference is \[t_2-t_1=t_3-t_2\]

OpenStudy (anonymous):

I'm really confused right now..

OpenStudy (astrophysics):

I would say so you jumped from a sequence from your last question to a series here, so just focus on the equation for now. The t here represents the term and the subscript is the address of the term, so \[t_1 = 3~~~t_2 = 1\] etc

OpenStudy (welshfella):

Common difference can be worked out by subtracting the first term from the second

OpenStudy (astrophysics):

\[t_2-t_1 = t_3-t_2 \implies 1-3 = -1-1 \implies -2 = -2 \checkmark\] so it works out, our common difference d = -2

OpenStudy (astrophysics):

With me so far?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

i'm just writing everything down in my notebook!

OpenStudy (astrophysics):

Ok cool :) Well now lets find a, what is a?

OpenStudy (anonymous):

is already one of the terms given, or do I have to find it?

OpenStudy (astrophysics):

a = first term in series

OpenStudy (anonymous):

A not already sorry

OpenStudy (anonymous):

so its 3

OpenStudy (astrophysics):

Good!

OpenStudy (astrophysics):

So we now have d = -2, a = 3, n = 11

OpenStudy (astrophysics):

Now we can use the equation I mentioned above \[S_n = \frac{ n }{ 2 }[2a+(n-a)d]\] can you plug it all in now

OpenStudy (anonymous):

sn=11/2 (2(3)+(11+3)-2) correct?

OpenStudy (astrophysics):

Close, your sign in (n-a) is off but it looks good \[S_{11} = \frac{ 11 }{ 2 }[2(3)+(11-3)(-2)]\]

OpenStudy (anonymous):

i got -55 ?

OpenStudy (astrophysics):

Yup, that looks good

OpenStudy (anonymous):

but it isn't one of the choices?

OpenStudy (astrophysics):

Mhm...

OpenStudy (astrophysics):

So is your question asking for \[S_{11}\]

OpenStudy (anonymous):

yes!

OpenStudy (astrophysics):

So it's exactly Find S11 for 3 + 1 + (-1) + (-3) +… right

OpenStudy (anonymous):

yes!

OpenStudy (astrophysics):

Mhm, everything looks right

OpenStudy (astrophysics):

Ok lets see...a = 3, d = -2, n = 11

OpenStudy (astrophysics):

I have located the problem it's the original equation I gave you, it should (n-1) not (n-a) sorry! \[S_n = \frac{ n }{ 2 }[2a+(n-1)d] = \frac{ 11 }{ 2 }[2(3)+(11-1)(-2)] = \]

OpenStudy (astrophysics):

Try now :)

OpenStudy (astrophysics):

So the equation is \[S_n = \frac{ n }{ 2 }[2a+(n-\color{red}1)d]\]

OpenStudy (anonymous):

I got 72.8? so -73?

OpenStudy (astrophysics):

Nope, you must've put in wrong

OpenStudy (astrophysics):

\[\frac{ 11 }{ 2 }[2(3)+(11-1)(-2)] = \frac{ 11 }{ 2 }[6+10(-2)] = \frac{ 11 }{ 2 }[6-20]=\frac{ 11 }{ 2 }(-14) = -77\]

OpenStudy (astrophysics):

Use your order of operations :-)